# Infrared-Assisted Single-Stage Framework for Joint Restoration and Fusion of Visible and Infrared Images under Hazy Conditions

### v2

@@ IV-D Ablation Study
We design six experimental settings to evaluate the effectiveness of each module. In the first setting, we remove the PGM-generated prompt ${\bm{\hat{P}}}_{ir}$, the haze density esitimation (HDE ) module, and the process of supplementing visible features with infrared features based on haze density, directly inputting ${\bm{F}}_{vi}$ into the Transformer Block for dehazing (denoted as “w/o ${\bm{F}}_{ir}$”) to assess the auxiliary role of infrared information. In the second setting, we replace the operation in Eq. (5) by directly adding ${\bm{\hat{F}}}_{ir}$ and ${\bm{F}}_{vi}$ (denoted as “w/o HDE”) to validate the impact of haze-density-based infrared integration. In the third setting, we remove the prompt embedding for ${\bm{\hat{P}}}_{ir}$ and the PEB module, injecting infrared features extracted by the encoder, guided by haze density ${\bm{H}}$, directly into ${\bm{F}}_{vi}$ (denoted as “w/o ${\bm{\hat{P}}}_{ir}$”) to evaluate the effect of prompt embedding. In the fourth setting, we test the effectiveness of the PEB module by removing it from IA-FRM, denoted as “w/o FR-PEB”. In the fifth setting, we remove ${\bm{\hat{P}}}_{vi}$ from MsPE-FM to verity the effectiveness of prompt embedding, denoted as “w/o ${\bm{\hat{P}}}_{vi}$”. In the sixth setting, we omit the PEB from the Fusion Block, performing feature concatenation and convolution directly (denoted as “w/o FB-PEB”) to validate its role in fusion.
@@
Table VII and Fig. 10 demonstrate the impact of each module on fusion performance. While removing any module leads to performance decrease, these changes may not be readily visible in the qualitative results but are clearly reflected in the quantitative data. As shown in Table VII, when infrared information is not used to supplement visible features, the metrics ${Q_{PI}}$ and ${Q_{NIQE}}$ increase significantly, and all fusion metrics decrease, confirming the effectiveness of incorporating infrared information to enhance visible features. If the HDE module is excluded, and infrared information is directly injected into the visible image, model performance does not improve. Instead, color distortion appears in the fusion results, and objective evaluation metrics decline to varying degrees. Similar issues are observed in the setting “w/o ${\bm{\hat{P}}}_{ir}$”. Additionally, for the settings “w/o FR-PEB”, ‘w/o ${\bm{\hat{P}}}_{vi}$” and “w/o FB-PEB”, the experimental results show slight degradation in detail retention, along with a decline in objective metrics, further validating the effectiveness of each module.
@@
TABLE VII: Quantitative results of six ablation experiments on the MSRS dataset, with the values in red indicating the best performance and those in blue indicating the second-best.
Models ${Q}_{PI}\downarrow$ ${Q}_{NIQE}\downarrow$ ${Q}_{AB/F}\uparrow$ ${Q}_{VIF}\uparrow$ ${Q}_{SCD}\uparrow$ ${Q}_{CV}\downarrow$ ${Q}_{MI}\uparrow$ ${Q}_{SF}\uparrow$ w/o $\bm{F}_{ir}$ 3.866 4.217 0.651 0.483 1.631 252.400 2.520 11.120 w/o HDE 3.887 4.319 0.644 0.464 1.585 253.014 2.337 10.657 w/o $\bm{\hat{P}}_{ir}$ 3.851 4.259 0.648 0.481 1.628 246.616 2.521 10.848 w/o FR-PEB 3.819 4.099 0.644 0.475 1.601 251.357 2.349 10.942 w/o $\bm{\hat{P}}_{vi}$ 3.820 4.212 0.638 0.478 1.598 294.239 2.132 10.902 w/o FB-PEB 3.846 4.115 0.645 0.481 1.656 241.963 2.607 10.854 Ours 3.811 4.110 0.652 0.490 1.662 238.420 2.720 11.050
IV-E Hyperparameters Analysis
@@
A Our method involves two key hyperparameters: $\alpha$, used to balance the $L1$-loss, and $M$, representing the number of Fusion Blocks in MsPE-FM. During training, they are set to 2 and 5, respectively. To validate the rationale behind these choices, we conduct a detailed analysis of their impact on the model’s performance.
@@
The impact of $\alpha$ on model performance. We fix $M$ at 5 and study the impact of different values of $\alpha$ within the range (0, 5] on the model performance. Table VIII presents the quantitative evaluation results of the fusion model on the MSRS dataset for various values of $\alpha$. The results indicate that when $\alpha$ is too small, the model performance does not reach its optimal level. Similarly, when $\alpha$ is too large, performance declines as well. The model achieves the best overall performance when $\alpha=2$, which validates the choice of $\alpha=2$ in this study.
The impact of $M$ on model performance. To analyze the impact of different values of $M$ on the model performance, we fix $\alpha$ at 2 and adjust $M$ within the range of [1, 7] to observe how the model performance changes. As shown in Table IX, the overall performance of the model improves with the number of Fusion Blocks. However, from the results for $M=5$ and $M=7$, it is evident that the performance gain has plateaued. Considering that increasing the number of Fusion Blocks significantly raises the model’s parameter count, we ultimately set $M$ to 5.
@@
TABLE VIII: Quantitative analysis of six different $\alpha$ models on the MSRS dataset under the condition of $M=5$. Red indicates the best, and blue indicates the second best.
$\alpha$ $Q_{MI}$uparrow $Q_{AB/F}$uparrow $Q_{CV}$downarrow $Q_{SCD}$uparrow $Q_{VIF}$uparrow $Q_{NIQE}$downarrow $Q_{PI}$downarrow $Q_{SF}$uparrow 0.5 2.711 0.650 241.934 1.646 0.489 4.211 3.816 11.115 1 2.764 0.660 239.645 1.656 0.483 4.209 3.821 11.083 2 2.720 0.652 238.420 1.662 0.490 4.110 3.811 11.050 3 2.743 0.658 243.385 1.650 0.488 4.105 3.819 11.049 4 2.787 0.656 240.587 1.657 0.489 4.149 3.838 11.054 5 2.030 0.372 387.993 1.572 0.345 5.321 5.401 7.059
@@
TABLE IX: Quantitative analysis of four different $M$ models on the MSRS dataset under the condition of $\alpha=2$. Red indicates the best, and blue indicates the second best.
$M$ $Q_{MI}$uparrow $Q_{AB/F}$uparrow $Q_{CV}$downarrow $Q_{SCD}$uparrow $Q_{VIF}$uparrow $Q_{NIQE}$downarrow $Q_{PI}$downarrow $Q_{SF}$uparrow 1 2.567 0.651 251.174 1.669 0.484 4.113 3.816 10.967 3 2.915 0.662 244.317 1.643 0.488 4.146 3.839 10.982 5 2.720 0.652 238.420 1.662 0.490 4.110 3.811 11.050 7 2.832 0.659 238.883 1.656 0.491 4.108 3.840 11.027
IV-F Complexity Analysis
@@
This paper employs a single-stage framework for hazy image fusion, which significantly reduces the model’s complexity and parameter count. To validate this advantage, we test the FLOPs and parameter count of each model and plot a bubble chart with the fusion metric $Q_{AB/F}$ to assess the model complexity. As shown in Fig. 11, our model achieves optimal performance while maintaining a low parameter count and the lowest computational complexity, fully demonstrating its efficiency and suitability for practical deployment.
@@
IV-G Limitations and Future Work

### v3

@@ 4.4 Ablation Study
We design seven experimental settings to evaluate the effectiveness of each module. In the first setting, we remove the PGM-generated prompt ${\bm{\hat{P}}}_{ir}$, the haze density esitimation (HDE ) module, and the process of supplementing visible features with infrared features based on haze density, directly inputting ${\bm{F}}_{vi}$ into the Transformer Block for dehazing (denoted as “w/o ${\bm{F}}_{ir}$”) to assess the auxiliary role of infrared information. In the second setting, we replace the operation in Eq. (9) by directly adding ${\bm{\hat{F}}}_{ir}$ and ${\bm{F}}_{vi}$ (denoted as “w/o HDE”) to validate the impact of haze-density-based infrared integration. In the third setting, we remove the prompt embedding for ${\bm{\hat{P}}}_{ir}$ and the PEB module, injecting infrared features extracted by the encoder, guided by haze density $\bm{H}$}, directly into ${\bm{F}}_{vi}$ (denoted as “w/o ${\bm{\hat{P}}}_{ir}$”) to evaluate the effect of prompt embedding. In the fourth setting, we test the effectiveness of the PEB module by removing it from IA-FRM, denoted as “w/o FR-PEB”. In the fifth setting, we remove ${\bm{\hat{P}}}_{vi}$ from MsPE-FM to verity the effectiveness of prompt embedding, denoted as “w/o ${\bm{\hat{P}}}_{vi}$”.In the sixth setting, we omit the PEB from the Fusion Block, performing feature concatenation and convolution directly (denoted as “w/o FB-PEB”) to validate its role in fusion. In the seventh setting, we remove the entire IA-FRM module (denoted as “w/o IA-FRM”) to evaluate the effectiveness of the dehazing component.
@@
Table 7 and Fig. 11 demonstrate the impact of each module on fusion performance. While removing any module leads to performance decrease, these changes may not be readily visible in the qualitative results but are clearly reflected in the quantitative data. As shown in Table 7, when infrared information is not used to supplement visible features, the metrics $Q_{PI}$} and $Q_{NIQE}$} increase significantly, and all fusion metrics decrease, confirming the effectiveness of incorporating infrared information to enhance visible features. If the HDE module is excluded, and infrared information is directly injected into the visible image, model performance does not improve. Instead, color distortion appears in the fusion results, and objective evaluation metrics decline to varying degrees. Similar issues are observed in the setting “w/o ${\bm{\hat{P}}}_{ir}$”. Additionally, for the settings “w/o FR-PEB”, ‘w/o ${\bm{\hat{P}}}_{vi}$” and “w/o FB-PEB”, the experimental results show slight degradation in detail retention, along with a decline in objective metrics, further validating the effectiveness of each module. In the “w/o IA-FRM” setting, when the IA-FRM module is removed and the hazy image pair is directly fed into the fusion network, it can be observed that the fusion results exhibit obvious color distortion and reduced contrast due to the presence of haze. Recognizable performance degradation is also reflected in the objective metrics, which further verifies the necessity of performing dehazing on visible images prior to fusion.
@@
Table 7: Quantitative results of seven ablation experiments on the MSRS dataset. The best and second-best performances are highlighted with Red and Blue backgrounds, respectively.
Models ${Q}_{PI}\downarrow$ ${Q}_{NIQE}\downarrow$ ${Q}_{AB/F}\uparrow$ ${Q}_{VIF}\uparrow$ ${Q}_{SCD}\uparrow$ ${Q}_{CV}\downarrow$ ${Q}_{MI}\uparrow$ ${Q}_{SF}\uparrow$ w/o $\bm{F}_{ir}$ 3.866 4.217 0.651 0.483 1.631 252.400 2.520 11.120 w/o HDE 3.887 4.319 0.644 0.464 1.585 253.014 2.337 10.657 w/o $\bm{\hat{P}}_{ir}$ 3.851 4.259 0.648 0.481 1.628 246.616 2.521 10.848 w/o FR-PEB 3.819 4.099 0.644 0.475 1.601 251.357 2.349 10.942 w/o $\bm{\hat{P}}_{vi}$ 3.820 4.212 0.638 0.478 1.598 294.239 2.132 10.902 w/o FB-PEB 3.846 4.115 0.645 0.481 1.656 241.963 2.607 10.854 w/o IA-FRM 4.241 4.584 0.552 0.439 1.332 310.824 1.967 9.183 Ours 3.811 4.110 0.652 0.490 1.662 238.420 2.720 11.050
4.5 Hyperparameters Analysis
@@
Our method involves three key hyperparameters: the coefficient $\alpha$ for balancing the $L1$ loss, the number of Transformer blocks $L$ in IA-FRM, and the number of Fusion Blocks $M$ in MsPE-FM. During training, these hyperparameters are set to 2, 5, and 5, respectively. To validate the rationale behind these choices, we further conduct a detailed analysis of their impact on the model’s performance.
@@
The impact of $\alpha$ on model performance. We fix $L$ and $M$ at 5 and study the impact of different values of $\alpha$ within the range (0, 5] on the model performance. Table 8 presents the quantitative evaluation results of the fusion model on the MSRS dataset for various values of $\alpha$. The results indicate that when $\alpha$ is too small, the model performance does not reach its optimal level. Similarly, when $\alpha$ is too large, performance declines as well. The model achieves the best overall performance when $\alpha=2$, which validates the choice of $\alpha=2$ in this study.
@@
Table 8: Quantitative analysis of six different $\alpha$ models on the MSRS dataset under the condition of $M=5$ and $L=5$. The best and second-best performances are highlighted with Red and Blue backgrounds, respectively.
$\alpha$ $Q_{MI}$ $\uparrow$ $Q_{AB/F}$ $\uparrow$ $Q_{CV}$ $\downarrow$ $Q_{SCD}$ $\uparrow$ $Q_{VIF}$ $\uparrow$ $Q_{NIQE}$ $\downarrow$ $Q_{PI}$ $\downarrow$ $Q_{SF}$ $\uparrow$ 0.5 2.711 0.650 241.934 1.646 0.489 4.211 3.816 11.115 1 2.764 0.660 239.645 1.656 0.483 4.209 3.821 11.083 2 2.720 0.652 238.420 1.662 0.490 4.110 3.811 11.050 3 2.743 0.658 243.385 1.650 0.488 4.105 3.819 11.049 4 2.787 0.656 240.587 1.657 0.489 4.149 3.838 11.054 5 2.030 0.372 387.993 1.572 0.345 5.321 5.401 7.059
The impact of $L$ on model performance. To analyze the impact of different values of $L$ on the model performance, we fix $\alpha$ at 2 and $M$ at 5, and vary $L$ within the range [1, 7] to observe the corresponding performance changes. As shown in Table 9, the overall performance of the model improves as the number of Transformer Blocks increases. However, from the results for $L=5$ and $L=7$, it is evident that the performance gain has already plateaued, with only marginal improvements. Considering the increase in model parameters associated with larger $L$, we ultimately set $L$ to 5.
@@
Table 9: Quantitative analysis of four different $L$ models on the MSRS dataset under the condition of $\alpha=2$ and $M=5$. The best and second-best performances are highlighted with Red and Blue backgrounds, respectively.
$L$ $Q_{MI}$ $\uparrow$ $Q_{AB/F}$ $\uparrow$ $Q_{CV}$ $\downarrow$ $Q_{SCD}$ $\uparrow$ $Q_{VIF}$ $\uparrow$ $Q_{NIQE}$ $\downarrow$ $Q_{PI}$ $\downarrow$ $Q_{SF}$ $\uparrow$ 1 2.353 0.592 273.282 1.492 0.411 4.526 4.113 9.988 3 2.592 0.636 259.921 1.589 0.468 4.391 3.981 10.147 5 2.720 0.652 238.420 1.662 0.490 4.110 3.811 11.050 7 2.769 0.659 240.167 1.641 0.493 4.082 3.826 10.992
The impact of $M$ on model performance. To analyze the impact of different values of $M$ on the model performance, we fix $\alpha$ at 2 and $L$ at 5, and vary $M$ within the range [1, 7] to observe the corresponding changes in performance. As shown in Table 10, the overall performance of the model improves as the number of Fusion Blocks increases. However, from the results for $M=5$ and $M=7$, it can be observed that the performance gain has already plateaued. Considering that further increasing the number of Fusion Blocks would significantly increase the number of model parameters, we ultimately set $M$ to 5.
Table 10: Quantitative analysis of four different $M$ models on the MSRS dataset under the condition of $\alpha=2$ and $L=5$. The best and second-best performances are highlighted with Red and Blue backgrounds, respectively.
$M$ $Q_{MI}$ $\uparrow$ $Q_{AB/F}$ $\uparrow$ $Q_{CV}$ $\downarrow$ $Q_{SCD}$ $\uparrow$ $Q_{VIF}$ $\uparrow$ $Q_{NIQE}$ $\downarrow$ $Q_{PI}$ $\downarrow$ $Q_{SF}$ $\uparrow$ 1 2.567 0.651 251.174 1.669 0.484 4.113 3.816 10.967 3 2.915 0.662 244.317 1.643 0.488 4.146 3.839 10.982 5 2.720 0.652 238.420 1.662 0.490 4.110 3.811 11.050 7 2.832 0.659 238.883 1.656 0.491 4.108 3.840 11.027
4.6 Complexity Analysis
@@
This paper employs a single-stage framework for hazy image fusion, which significantly reduces the model’s complexity and parameter count. To validate this advantage, we test the FLOPs and parameter count of each model and plot a bubble chart, where the bubble size denotes the fusion metric $Q_{AB/F}$ computed on the MSRS dataset. As shown in Fig. 12, our model achieves optimal performance while maintaining a low parameter count and the lowest computational complexity, fully demonstrating its efficiency and suitability for practical deployment.
@@
4.7 Limitations and Future Work
@@ 5 Conclusion
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants (62161015, 62276120), and in part by the Yunnan Fundamental Research Projects under Grants (202301AV070004, 202501AS070123, 202401AS070640).